# TwinBrainVLA[Uncaptioned image]: Unleashing the Potential of Generalist VLMs for Embodied Tasks via Asymmetric Mixture-of-Transformers

### v1

@@ Abstract
Standard Vision-Language-Action (VLA) models typically fine-tune a monolithic Vision-Language Model (VLM) backbone explicitly for robotic control. However, this approach creates a critical tension between maintaining high-level general semantic understanding and learning low-level, fine-grained sensorimotor skills, often leading to "catastrophic forgetting" of the model’s open-world capabilities. To resolve this conflict, we introduce TwinBrainVLA[Uncaptioned image], a novel architecture that coordinates a generalist VLM retaining universal semantic understanding and a specialist VLM dedicated to embodied proprioception for joint robotic control. TwinBrainVLA synergizes a frozen "Left Brain", which retains robust general visual reasoning, with a trainable "Right Brain", specialized for embodied perception, via a novel Asymmetric Mixture-of-Transformers (AsyMoT) mechanism. This design allows the Right Brain to dynamically query semantic knowledge from the frozen Left Brain and fuse it with proprioceptive states, providing rich conditioning for a Flow-Matching Action Expert to generate precise continuous controls. Extensive experiments on SimplerEnv and RoboCasa benchmarks demonstrate that TwinBrainVLA achieves superior manipulation performance compared to state-of-the-art baselines while explicitly preserving the comprehensive visual understanding capabilities of the pre-trained VLM, offering a promising direction for building general-purpose robots that simultaneously achieve high-level semantic understanding and low-level physical dexterity.
TwinBrainVLA[Uncaptioned image]: Unleashing the Potential of Generalist VLMs for Embodied Tasks via Asymmetric Mixture-of-Transformers
@@ 1 Introduction
The pursuit of embodied artificial intelligence has recently converged on the paradigm of Vision-Language-Action (VLA) models (Kim et al., 2024; Intelligence et al., 2025; NVIDIA et al., 2025). By grafting robotic control heads onto pre-trained Vision-Language Models (VLMs), these systems inherit the rich semantic reasoning and open-world generalization capabilities of large-scale internet data. This unified approach promises a robot that can not only recognize a "coffee mug" but also understand the instruction "place the mug on the coaster" and execute the corresponding motor commands.
However, a fundamental tension lies at the heart of current VLA architectures: the misalignment between the optimization objectives of semantic pre-training and robotic fine-tuning. Pre-trained VLMs are originally optimized for general semantic understanding and dialogue. In contrast, during VLA fine-tuning, the backbone is repurposed to provide precise, spatially-grounded conditioning for the Action Expert to enable precise robotic control. Forcing a single VLM backbone to satisfy these divergent objectives simultaneously inevitably precipitates "catastrophic forgetting"(Hancock et al., 2025b; Zhang et al., 2026; Zhou et al., 2025; Fang et al., 2025), where the model sacrifices its general linguistic capabilities to adapt to the specialized demands of robotic actuation. More critically, this catastrophic forgetting undermines the fundamental premise of the VLA paradigm, which aims to leverage the general world understanding acquired during VLM pre-training to enable generalized robotic control. However, current fine-tuning methodologies paradoxically strip the VLM of this very capability.
To resolve this conflict, we draw inspiration from the biological principle of hemispheric lateralization, where the human brain allocates distinct cognitive functions to specialized hemispheres—logic and language to the left, spatial attention and motor coordination to the right. We propose that VLA models should mimic this decoupling. By separating general semantic understanding from embodied action estimation, we cY-an achieve a system that is both cognitively capable and physically skillful.
@@
In this paper, we introduce TwinBrainVLA[Uncaptioned image], a novel VLA framework that orchestrates two isomorphic VLM pathways via an asymmetric dual-stream joint training strategy. Structurally, our model consists of a "Left Brain" and a "Right Brain". The Left Brain is a frozen, pre-trained VLM that retains robust open-world knowledge and instruction following capabilities. The Right Brain, initialized with the same architecture, is fully trainable and specialized for embodied perception. These two streams are seamlessly fused through a Asymmetric Mixture-of-Transformers (AsyMoT) mechanism, which allows for joint attention over hidden states without sharing parameters. Crucially, the specialized representations from the Right Brain serve as the primary condition for our action expert utilizing flow-matching for precise action generation. This design ensures that the action expert receives spatially rich, task-aligned embeddings from the Right Brain, while the Left Brain explicitly preserves the model’s general semantic understanding capabilities.
Our contributions are threefold:
  • We propose TwinBrainVLA[Uncaptioned image], the first VLA architecture to explicitly decouple general semantic understanding and embodied perception via an asymmetric dual-stream design, resolving the training conflict inherent in single-backbone VLAs.
  • We introduce the Asymmetric Mixture-of-Transformers (AsyMoT) mechanism for the information interaction of two isomorphic VLM pathways to build the VLA backbone, and employ an asymmetric parameter freezing strategy to achieve the joint training of the dual models.
  • Extensive comparative experiments and evaluations on the SimplerEnv and RoboCasa benchmarks, demonstrates the effectiveness of the TwinBrainVLA architecture, the AsyMoT mechanism, and the proposed training strategy.
@@ 2 Related Work
Multimodal understanding. The landscape of computer vision and natural language processing has been revolutionized by Vision Language Models (VLMs) (Liu et al., 2023; Bai et al., 2025b, a). By seamlessly integrating visual encoders with Large Language Models (LLMs) through sophisticated projection layers or adapter modules (Jiang et al., 2025a), these models exhibit emergent abilities in holistic scene understanding, visual question answering (VQA), and complex math reasoning. Given the inherent limitations of general-purpose VLMs in spatial perception (Xiao et al., 2026; Dang et al., 2025; Zhang et al., 2025b; Chen et al., 2025; Yang et al., 2025c), recent studies have increasingly focused on employing post-training strategies to enhance spatial intelligence (Yang et al., 2025a; Tan et al., 2025; Zhou et al., 2026; Zhu et al., 2025; Li et al., 2025) and construct specialized embodied foundation models tailored for embodied scenarios (Team et al., 2025a; Yang et al., 2025a; Hao et al., 2025; Lin et al., 2025c).
@@
Vision-Language-Action (VLA) Models. Building upon the zero-shot generalization and rich semantic priors of VLMs, VLA models have emerged as a scalable paradigm for embodied intelligence Pertsch et al. (2025); Zheng et al. (2025a); Wen et al. (2025); Shukor et al. (2025); Zheng et al. (2025b); Lin et al. (2025b); Cai et al. (2026). By fine-tuning pre-trained VLMs on large-scale robotic manipulation datasets, these models learn to map multimodal instructions and observations to low-level control signals (Jiang et al., 2025b; Fan et al., 2025; Zhai et al., 2025). To mitigate the degradation of general conversational skills during VLA training, recent approaches have integrated general visual-dialogue datasets (Mazzaglia et al., 2025; Yang et al., 2025b; Hancock et al., 2025a; Zhou et al., 2025) and reasoning mechanisms into the VLA framework (Zawalski et al., 2025; Yin et al., 2025; Lin et al., 2025a; Qu et al., 2025a; Ye et al., 2025; Huang et al., 2025; Guo et al., 2025; Lee et al., 2025; Huang et al., 2026).
3 Motivation: Catastrophic Forgetting in VLA Models
Standard Vision-Language-Action (VLA) models typically adopt a transfer learning paradigm. Formally, let a pre-trained Vision-Language Model (VLM) be parameterized by $\theta_{\text{pre}}$, which has been aligned on massive internet-scale image-text pairs to maximize the likelihood of generating semantic text responses $y$ given an image $v$ and instruction $x$. To adapt this backbone for robotic manipulation, the model is fine-tuned on a robotic demonstration dataset $\mathcal{D}_{\text{robot}}=\{(v,x,a)_{i}\}_{i=1}^{N}$, where $a$ represents the robot’s action (e.g., end-effector pose or joint angles). The training objective shifts to minimizing the action prediction loss $\mathcal{L}_{\text{action}}$:
$\theta^{*}=\arg\min_{\theta}\sum_{(v,x,a)\in\mathcal{D}_{\text{robot}}}-\log P(a|v,x;\theta),$ (1)
initialized with $\theta=\theta_{\text{pre}}$.
While this process effectively imparts manipulation skills—enabling the model to control servos and execute tasks—it introduces a critical side effect known as catastrophic forgetting. The distribution of $\mathcal{D}_{\text{robot}}$ differs fundamentally from the diverse visual-linguistic data seen during pre-training (Kachaev et al., 2025; Hancock et al., 2025a; Zhang et al., 2025a). As the parameters $\theta$ are updated to minimize the high-frequency spatial error in action space, the delicate semantic alignment established in $\theta_{\text{pre}}$ is often overwritten. Consequently, the resulting VLA model suffers from a severe degradation in its original capabilities: it loses the ability to follow complex open-ended instructions, fails to describe visual scenes accurately, and exhibits weakened general semantic reasoning. This phenomenon turns the VLM backbone from a "generalist" into a "specialist", sacrificing its linguistic brain to gain a robotic body.
@@ 4 Method: TwinBrainVLA[Uncaptioned image]
In this section, we present TwinBrainVLA[Uncaptioned image], a unified VLA framework designed to disentangle high-level semantic reasoning from fine-grained sensorimotor control. As illustrated in Figure 2, our architecture operates on an asymmetric dual-stream paradigm to resolve the "catastrophic forgetting" dilemma. The framework comprises three core components: (1) a frozen "Left Brain" that preserves open-world visual-linguistic knowledge; (2) a trainable "Right Brain" that specializes in embodied motor control; and (3) a Flow-Matching Action Expert that generates precise continuous actions. The two brains interact via a novel Asymmetric Mixture-of-Transformers (AsyMoT) mechanism, enabling the specialist to leverage generalist semantics without corrupting them. We detail the Dual-VLM architecture and AsyMoT fusion strategy in Sec. 4.1, describe the flow-matching policy formulation in Sec. 4.2, and finally present our joint-training objective in Sec. 4.3.
@@
4.1 Asymmetric Dual-VLM as Backbone
The backbone of TwinBrainVLA consists of two parallel Vision-Language Models (VLMs), denoted as the "Left Brain" ($\mathcal{M}_{L}$) and the "Right Brain" ($\mathcal{M}_{R}$). Both streams are initialized with identical pre-trained weights (e.g., Qwen3-VL series (Bai et al., 2025a)) but serve distinct roles during training.
Dual-Stream Inputs and State Encoding. The input to the system comprises a visual observation $\mathbf{I}$, a natural language instruction $\mathbf{T}$, and the robot’s proprioceptive state $\mathbf{s}\in\mathbb{R}^{d_{s}}$ (e.g., joint angles and end-effector pose). The Left Brain functions as a pure semantic generalist. It receives only the vision and language tokens:
$\mathbf{H}_{L}^{0}=[\mathcal{V}(\mathbf{I});\mathcal{T}(\mathbf{T})],$ (2)
where $\mathcal{V}$ and $\mathcal{T}$ represent the vision encoder and text tokenizer, respectively.
In contrast, the Right Brain is responsible for embodied control and must be grounded in the robot’s physical state. To achieve this, we introduce a lightweight State Encoder $\phi$, modeled as a Multi-Layer Perceptron (MLP), which projects the low-level proprioceptive state $\mathbf{s}$ into the VLM’s embedding space. These state tokens are appended to the multimodal input sequence:
$\mathbf{H}_{R}^{0}=[\mathcal{V}(\mathbf{I});\mathcal{T}(\mathbf{T});\phi(\mathbf{s})].$ (3)
This design ensures that the Right Brain explicitly attends to the robot’s current configuration, a critical requirement for closed-loop control.
Asymmetric Mixture-of-Transformers (AsyMoT). The core innovation of our backbone lies in how these two streams interact. We propose the AsyMoT mechanism to fuse general semantic knowledge into the robotic control stream. Let $\mathbf{H}_{L}^{l}$ and $\mathbf{H}_{R}^{l}$ denote the hidden states of the Left and Right Brains at layer $l$, respectively.
@@ 4.1 Asymmetric Dual-VLM as Backbone
This asymmetric flow ensures a strict hierarchy: the Left Brain acts as a stable "semantic anchor" providing high-level reasoning features, while the Right Brain dynamically fuses these semantics with fine-grained proprioceptive cues to reason about spatial actions. The final hidden states of the Right Brain, $\mathbf{H}_{R}^{final}$, are then passed to the Action Expert.

### v2

@@ Abstract
The fundamental premise of Vision-Language-Action (VLA) models is to harness the extensive general capabilities of pre-trained Vision-Language Models (VLMs) for generalized embodied intelligence. However, standard robotic fine-tuning inevitably disrupts the pre-trained feature space, leading to "catastrophic forgetting" that compromises the general visual understanding we aim to leverage. To effectively utilize the uncorrupted general capabilities of VLMs for robotic tasks, we propose TwinBrainVLA[Uncaptioned image], which coordinates two isomorphic VLM pathways: a frozen generalist (also called "Left Brain") and a trainable specialist (also called "Right Brain"). Our architecture utilizes a Asymmetric Mixture-of-Transformers (AsyMoT) mechanism, enabling the Right Brain to dynamically query and fuse intact semantic knowledge from the Left Brain with proprioceptive states. This fused representation conditions a flow-matching action expert for precise continuous control. Empirical results on SimplerEnv and RoboCasa benchmarks demonstrate that by explicitly retaining general capabilities, TwinBrainVLA achieves substantial performance gains over baseline models in complex manipulation tasks.
TwinBrainVLA[Uncaptioned image]: Unleashing the Potential of Generalist VLMs
for Embodied Tasks via Asymmetric Mixture-of-Transformers
@@ 1 Introduction
The pursuit of embodied artificial intelligence has recently converged on the paradigm of Vision-Language-Action (VLA) model (Kim et al., 2024; Intelligence et al., 2025; NVIDIA et al., 2025). By grafting robotic control heads onto pre-trained Vision-Language Models (VLMs), the fundamental premise is to harness the rich semantic reasoning and open-world generalization capabilities learned from internet-scale data to enable generalized robotic control. Ideally, such a system should not merely execute motor commands but leverage its "VLM brain" to understand complex instructions, reason about unseen objects, and plan over long horizons.
However, a critical paradox undermines current VLA architectures: the optimization process required to learn low-level sensorimotor control inevitably disrupts the high-level semantic features we aim to leverage. During standard robotic fine-tuning, the VLM backbone is aggressively updated to minimize action prediction errors on narrow, domain-specific datasets. This creates a conflict between the original objective of semantic understanding and the new objective of proprioceptive control, precipitating "catastrophic forgetting(Hancock et al., 2025b; Zhang et al., 2026; Zhou et al., 2025; Fang et al., 2025). Consequently, the model sacrifices its pre-trained general capabilities to become a specialized controller. This defeats the original purpose of the VLA paradigm: instead of inheriting a generalized "brain," the fine-tuned model degenerates into a localized policy, stripping away the very open-world understanding required for robust generalization.
Therefore, the central challenge is to effectively utilize the uncorrupted general capabilities of VLMs to enhance robotic performance. We posit that a generalizable VLA model must simultaneously possess both general visual understanding capabilities and low-level servo control proficiency. Drawing inspiration from the biological principle of hemispheric lateralization, where the brain allocates distinct functions to specialized hemispheres, we propose that a VLA should maintain a dedicated "semantic anchor" separate from its "motor controller."
@@
In this paper, we introduce TwinBrainVLA[Uncaptioned image], a novel VLA framework that orchestrates two isomorphic VLM pathways via an asymmetric dual-stream joint training strategy. Structurally, our model comprises a frozen "Left Brain" (generalist) and a trainable "Right Brain" (specialist). The Left Brain preserves the intact, open-world knowledge of the pre-trained VLM, while the Right Brain specializes in processing proprioceptive states and generating actions. Crucially, we propose a novel Asymmetric Mixture-of-Transformers (AsyMoT) mechanism, which enables the trainable Right Brain to dynamically query and fuse semantic features from the frozen Left Brain. This design ensures that the action expert is conditioned on uncorrupted general knowledge, allowing the robot to "think" with a generalist brain while "acting" with a specialist body.
Our contributions are summarized as follows:
  • We quantitatively analyzed the impairment of the general capabilities of VLMs caused by VLA training through preliminary experiments, namely the occurrence of the "catastrophic forgetting" phenomenon.
  • We propose TwinBrainVLA[Uncaptioned image], an asymmetric dual-stream architecture that structurally decouples semantic understanding from embodied control, allowing the system to leverage pre-trained capabilities without disruption.
  • We introduce the Asymmetric Mixture-of-Transformers (AsyMoT) mechanism to facilitate efficient information interaction between the two VLM pathways, enabling the specialist Right Brain to dynamically query general knowledge from the frozen Left Brain.
  • Extensive experiments and evaluations on the SimplerEnv and RoboCasa benchmarks, as well as real-robot settings, demonstrate the effectiveness of the TwinBrainVLA architecture, the AsyMoT mechanism, and our proposed training strategy.
@@ 2 Related Work
Multimodal understanding. The landscape of computer vision and natural language processing has been revolutionized by VLMs (Liu et al., 2023b; Bai et al., 2025b, a). By seamlessly integrating visual encoders with Large Language Models (LLMs) through sophisticated projection layers or adapter modules (Jiang et al., 2025a), these models exhibit emergent abilities in semantic understanding and visual question answering (VQA). Given the inherent limitations of general-purpose VLMs in spatial perception (Xiao et al., 2026; Dang et al., 2025; Zhang et al., 2025; Chen et al., 2025; Yang et al., 2025b), recent studies have increasingly focused on employing post-training strategies to enhance spatial intelligence (Yang et al., 2025a; Tan et al., 2025; Zhou et al., 2026; Zhu et al., 2025; Li et al., 2025) and construct specialized embodied foundation models tailored for embodied scenarios (Team et al., 2025a; Yang et al., 2025a; Hao et al., 2025; Lin et al., 2025c).
@@
Vision-Language-Action (VLA) Models. Building upon the zero-shot generalization and rich semantic priors of VLMs, VLA models have emerged as a scalable paradigm for embodied intelligence Pertsch et al. (2025); Zheng et al. (2025a); Wen et al. (2025); Shukor et al. (2025); Zheng et al. (2025b); Lin et al. (2025b); Cai et al. (2026). By fine-tuning pre-trained VLMs on large-scale robotic manipulation datasets, these models learn to map multimodal instructions and observations to low-level control signals (Jiang et al., 2025b; Fan et al., 2025; Zhai et al., 2025). To mitigate the degradation of general conversational capabilities in VLMs during VLA training, several approaches have been proposed. LangForce (Lian et al., 2026) attempts to enhance VLA performance by introducing mechanisms to mitigate visual shortcuts. ChatVLA (Zhou et al., 2025) and VLM2VLA (Hancock et al., 2025a) attempts to incorporate general dialogue data for co-training; however, this method still suffers from catastrophic forgetting and requires careful curation of general QA data. ThinkAct (Huang et al., 2025) and related works (Yin et al., 2025; Lin et al., 2025a; Lee et al., 2025) introduce Chain-of-Thought (CoT) reasoning into VLA training. Nevertheless, current mainstream VLA datasets lack CoT supervision signals, and the annotation process remains expensive and difficult to scale. Instead of relying on extensive data engineering, we propose a structural solution that explicitly decouples semantic preservation from motor learning, enabling the VLA to directly harness the uncorrupted general capabilities of the VLM backbone.
Figure 2: Empirical Evidence of Catastrophic Forgetting. (a) Overview of two prevailing training paradigms: standard VLA-Training (utilizing robot-only data) and Co-Training (mixing robot data with general vision-language data). (b) Evaluation of the VLM’s general visual understanding capabilities across four benchmarks (POPE (Li et al., 2023b), SEED-Bench (Li et al., 2023a), ScienceQA (Lu et al., 2022), and MME (Fu et al., 2025)) during the VLA fine-tuning process. Both strategies exhibit a precipitous decline in performance as training steps increase, demonstrating that co-training with general data is insufficient to prevent the collapse of the VLM’s pre-trained semantic reasoning when optimizing for robotic control.
3 Motivation: The Dilemma of Utilizing General Capabilities in VLA
3.1 The Conflict Between Specialization and Generalization
Standard VLA models are built on a compelling premise: by fine-tuning pre-trained VLMs on robotic datasets, we aim to endow robots with the rich semantic reasoning and open-world understanding inherent in the backbone. Ideally, the VLM should serve as a "generalist brain" that guides the robotic body through complex, unseen environments.
However, a fundamental paradox exists in the current transfer learning paradigm. To adapt the VLM for precise sensorimotor control, the model is fine-tuned on robotic datasets ($\mathcal{D}_{robot}$) that are fundamentally different from the massive internet-scale data used during pre-training. This aggressive optimization for low-level action prediction inevitably disrupts the pre-trained feature space, triggering "catastrophic forgetting". This phenomenon turns the VLM backbone from a "generalist" into a "specialist", sacrificing its linguistic brain to gain a robotic body.
3.2 Empirical Analysis: The Severity of Catastrophic Forgetting
To empirically quantify the extent of catastrophic forgetting discussed above, we conducted a preliminary study using VLM backbones: Qwen2.5-VL-3B-Instruct and Qwen3-VL-4B-Instruct. We evaluate the models’ general visual capabilities before and after fine-tuning under two distinct paradigms: (1) VLA Training: The model is fine-tuned exclusively on robotic trajectory data to optimize action generation. (2) Co-Training: Following recent strategies (Mazzaglia et al., 2025; Zhou et al., 2025) to mitigate forgetting, we mix general visual-QA data with robotic data at a 1:1 ratio during fine-tuning, aiming to preserve visual conversational capabilities.
Result Analysis. The results are presented in Figure 2b. First, standard VLA Training leads to a complete collapse of general visual understanding. For instance, the POPE score of Qwen3-VL drops from 88.87% to near-zero (0.04%). This confirms that optimizing for proprioceptive action control aggressively overwrites the semantic features required for VQA tasks. More critically, the co-training strategy also fails to prevent this degradation.
Key Findings 1. Training VLA models with robotics data induces severe catastrophic forgetting in the VLM, which undermines the original premise of the paradigm that builds VLAs from VLMs. 2. While Co-Training is frequently suggested as a countermeasure, our experiments show that it merely mitigates the symptoms rather than providing a fundamental solution.
These findings expose a critical bottleneck: standard fine-tuning effectively erases the ’VLM brain’ before it can be deployed for robotic reasoning, rendering it impossible to harness pre-trained general knowledge for complex, open-ended tasks. While the success of multimodal applications has proven that general understanding is essential for downstream performance which is also a key insight driving the VLA paradigm. However, current training paradigms lead to severe catastrophic forgetting, which directly contradicts the initial design philosophy of VLAs. Therefore, the challenge remains: How can we construct VLA models to leverage the uncorrupted multimodal understanding capabilities of VLMs to enhance their own performance?
@@ 4 Method: TwinBrainVLA[Uncaptioned image]
In this section, we present TwinBrainVLA[Uncaptioned image], a novel framework designed to enable VLA models to effectively utilize the general visual understanding capabilities of pre-trained VLMs for embodied tasks. While standard monolithic architectures suffer from feature degradation during fine-tuning, our approach aims to retain the full spectrum of pre-trained semantic knowledge and make it explicitly accessible for robotic control. To this end, we introduce an asymmetric dual-stream design that structurally separates the preservation of general capabilities from the learning of sensorimotor skills, ensuring that the control policy can leverage uncorrupted multimodal understanding capabilities.
@@
4.1 Asymmetric Dual-VLM Backbone
The backbone of TwinBrainVLA comprises two parallel VLM pathways: the "Left Brain" ($\mathcal{M}_{L}$) and the "Right Brain" ($\mathcal{M}_{R}$). Both streams are initialized with identical pre-trained weights (e.g., Qwen3-VL series) but play distinct roles to facilitate the utilization of general capabilities for robot control.
The "Left Brain" as a General Capability Reservoir. The Left Brain serves as the provider of general visual understanding. To ensure that the pre-trained capabilities remain intact, it is kept strictly frozen throughout the training phase. It processes the visual observation $I$ and instruction $T$ to generate semantic representations:
$H_{L}^{0}=[\mathcal{V}(I);\mathcal{T}(T)]$ (1)
where $\mathcal{V}$ and $\mathcal{T}$ denote the vision encoder and text tokenizer. By freezing this pathway, the Left Brain acts as a stable semantic anchor, offering a continuous source of open-world knowledge that is immune to the catastrophic forgetting typically induced by robotic fine-tuning.
The "Right Brain" as a Semantic-Augmented Controller. The Right Brain is a trainable specialist responsible for executing robotic actions. To ground the model in the physical world, we introduce a lightweight State Encoder $\phi$, which projects the robot’s proprioceptive state $s$ (e.g., joint angles) into the VLM’s embedding space:
$H_{R}^{0}=[\mathcal{V}(I);\mathcal{T}(T);\phi(s)]$ (2)
Unlike a standard VLA that relies solely on its own degrading parameters, the Right Brain is designed to actively harness the capabilities of the Left Brain to inform its decision-making.
Asymmetric Mixture-of-Transformers (AsyMoT). Mixture-of-Transformers (MoT) (Liang et al., 2025) is a model architecture that implements forward propagation by connecting two independent Transformer networks via joint attention at each corresponding layer. Its advantage lies in the ability to facilitate information interaction between multiple modalities while still maintaining their independent computation. Inspired by this work, we propose the Asymmetric Mixture-of-Transformers (AsyMoT) mechanism to enable the Right Brain to utilize the general visual understanding from the Left Brain. This mechanism establishes a unidirectional information bridge, allowing the control policy to query the frozen general knowledge.
Let $\mathbf{H}_{L}^{l}$ and $\mathbf{H}_{R}^{l}$ denote the hidden states of the Left and Right Brains at layer $l$, respectively.
@@ 4.1 Asymmetric Dual-VLM Backbone
AsyMoT v.s. Cross-Attention. The key to the AsyMoT mechanism lies in the asymmetric joint attention at each layer. It allows the ’right brain’ to attend to both the ’left brain’ and itself, whereas the ’left brain’ is restricted to attending only to itself, resulting in an asymmetric phenomenon. In contrast, Cross Attention is a different multimodal fusion architecture. It enables a Query from one modality to attend to the Key-Values of another modality, but precludes it from attending to itself. This constitutes the critical distinction between AsyMoT and Cross Attention."